# A refined Weyl character formula for comodules on $\operatorname{GL}(2,A)$

### v2

@@ Abstract.
Let $A$ be any commutative unital ring and let $\operatorname{GL}(2,A)$ be the general linear group on $A$ of rank $2$. We study the representation theory of $\operatorname{GL}(2,A)$ and the symmetric powers $\operatorname{Sym}^{d}(V)$, where $(V,\Delta)$ is the standard right comodule on $\operatorname{GL}(2,A)$. We prove a refined Weyl character formula for $\operatorname{Sym}^{d}(V)$. There is for any integer $d\geq 1$ a (canonical) refined weight space decomposition $\operatorname{Sym}^{d}(V)\cong\oplus_{i}\operatorname{Sym}^{d}(V)^{i}$ where each direct summand $\operatorname{Sym}^{d}(V)^{i}$ is a comodule on $N\subseteq\operatorname{GL}(2,A)$. Here $N$ is the schematic normalizer of the diagonal torus $T\subseteq\operatorname{GL}(2,A)$. We prove a character formula for the direct summands of $\operatorname{Sym}^{d}(V)$ for any integer $d\geq 1$. This refined Weyl character formula implies the classical Weyl character formula. As a Corollary we get a refined Weyl character formula for the pull back $\operatorname{Sym}^{d}(V\otimes K)$ as a comodule on $\operatorname{GL}(2,K)$ where $K$ is any field. We also calculate explicit examples involving the symmetric powers, symmetric tensors and their duals.
@@ 1. Introduction
In a recent paper (see [9]) I studied the notion ”good filtration” for torsion free comodules on $\operatorname{SL}(2,\mathbf{Z})$ and I proved the existence of a ”virtual Clebsch-Gordan formula” for the tensor product $\operatorname{Sym}^{m}(V)\otimes\operatorname{Sym}^{n}(V)$ of the symmetric powers of the standard right comodule $(V,\Delta)$ on the group scheme $\operatorname{SL}(2,A)$ where $A$ is an arbitrary commutative unital ring and $d\geq 1$ is an integer. This gives an infinite set of non trivial examples of torsion free finite rank comodules on $\operatorname{SL}(2,\mathbf{Z})$ equipped with a good filtration. In general one wants to give a definition of the notion ”good filtration” and to classify torsion free finite rank comodules with a good filtration. In [9] I found examples of non isomorphic comodules on $\operatorname{SL}(2,\mathbf{Z})$ that become isomorphic when we ”restrict to the generic fiber”. There is a canonical morphism of group schemes $\pi:\operatorname{SL}(2,\mathbf{Q})\rightarrow\operatorname{SL}(2,\mathbf{Z})$ and I proved existence of non isomorphic finite rank torsion free comodules $W_{1},W_{2}$ with $\pi^{*}W_{1}\cong\pi^{*}W_{2}$ an isomorphism of comodules on $\operatorname{SL}(2,\mathbf{Q})$. In order to define the notion ”good filtration” for comodules on $\operatorname{SL}(2,\mathbf{Z})$, we must for each integer $d\geq 1$ classify comodules $W$ on $\operatorname{SL}(2,\mathbf{Z})$ with the property that the pull back to $\operatorname{SL}(2,\mathbf{Q})$ becomes isomorphic to $\operatorname{Sym}^{d}(V)$ where $V:=\mathbf{Q}^{2}$ is the ”standard right comodule” of $\operatorname{SL}(2,\mathbf{Q})$. There is by [9] no highest weight theory or ”complete reducibility property” for torsion free comodules on $\operatorname{SL}(2,\mathbf{Z})$ and $\operatorname{GL}(2,\mathbf{Z})$, and for this reason we need new invariants and new methods in the study of such comodules and the aim of this paper is to introduce a new structure on a class of torsion free comodules on $\operatorname{SL}(2,\mathbf{Z})$.
In this paper I prove the following: Let $A$ be any commutative unital ring and let $G:=\operatorname{GL}(2,A)$ or $\operatorname{SL}(2,A)$. I define the schematic normalizer $N\subseteq G$, which is a closed subgroup scheme of $G$ containing the diagonal torus $T$ as a closed normal subgroup scheme. The schematic normalizer is a strict closed subgroup scheme of the functor normalizer $N_{G}(T)$ as defined in [8]. Let $\sigma,q:N\times T\rightarrow N$ be the action and projection map.
@@
For any map $f:N\rightarrow Z$ to an affine scheme $Z$ with $f\circ\sigma=f\circ q$ there is a unique map $f^{*}:N/T\rightarrow Z$ with $f^{*}\circ\pi=f$. Hence the affine group scheme $N/T$ is uniquely determined up to isomorphism.
@@
Then I prove the existence of a refined weight space decomposition of any symmetric power $W:=\operatorname{Sym}^{d}(V)$ where $d\geq 1$ is an integer and where $(V,\Delta)$ is the standard right comodule on $\operatorname{GL}(2,A)$. There is a canonical decomposition $W\cong\oplus_{\lambda\in\Lambda}W_{\lambda}$ where each summand $W_{\lambda}$ is a comodule on the schematic normalizer $N\subseteq\operatorname{GL}(2,A)$ of the diagonal torus $T\subseteq\operatorname{GL}(2,A)$. Then I prove a refined Weyl character formula for each summand $W_{\lambda}$ of $W$. The refined character formula implies the classical Weyl character formula (this is Theorem 6.4 in the paper):
@@
There is a minor mistake in [8] on the properties of $N_{G}(T)$ and $N_{G}(T)/T$. It is not true that there is an isomorphism $N_{G}(T)/T\cong\operatorname{Spec}(k[W_{G}])$ where $W_{G}$ is the Weyl group of $G$ and $\operatorname{Spec}(k[W_{G}])$ is the finite group scheme associated to $W_{G}$. This follows from an explicit calculation of the Hopf algebra $k[N/T]$ given in this paper. This will not be a problem in this paper, since we will calculate the roots of $\operatorname{GL}(2,A)$ and $\operatorname{SL}(2,A)$ explicitly using the coordinate ring $k[G]$ and we will not use results related to the relation between the Weyl group and the functor normalizer $N_{G}(T)$ from [8]. We will also calculate the comodule structure of any comodule explicitly using coordinate rings. We will not use the functor of points when working with $N$ and the Weyl group.
As an application of the methods introduced in the paper I prove the following Theorems (see Theorem 7.3). Let $G:=\operatorname{GL}(2,A)$ where $A$ is any commutative ring and let $(V,\Delta)$ be the standard left comodule on $G$. Let $\operatorname{Sym}^{2}(V),\operatorname{S}_{2}(V),\operatorname{Sym}^{2}(V)^{*}$ and $\operatorname{S}_{2}(V)^{*}$ be the symmetric power, the symmetric tensor and the dual of the symmetric power and symmetric tensor of $V$. The symmetric tensor is the invariants $\operatorname{S}_{2}(V)\subseteq V\otimes_{A}V$ under the canonical action of the symmetric group on two elements.
@@
Let $G:=\operatorname{GL}(2,A)$ with $A$ any commutative ring, and let $N\subseteq G$ be the schematic normalizer of the diagonal torus. There is an isomorphism of $k[N]$-comodules (and $k[T]$-comodules) $\operatorname{Sym}^{2}(V)\cong\operatorname{S}_{2}(V)$. Let $G:=\operatorname{GL}(2,\mathbf{Z})$ with $\mathbf{Z}$ the ring of integers. It follows the $k[G]$-comodules $\operatorname{Sym}^{2}(V)$ and $\operatorname{S}_{2}(V)$ are not isomorphic.
@@
It is hoped the results in this paper may be interesting in the study of torsion free comodules on $\operatorname{SL}(n,A)$ and $\operatorname{GL}(n,A)$ in the future. One problem one wants to study is the problem of classifying $\operatorname{SL}(2,\mathbf{Z})$ modules $W$ whose base change to $\operatorname{SL}(2,\mathbf{Q})$ are isomorphic to $\operatorname{Sym}^{d}(\mathbf{Q}^{2})$ where $\mathbf{Q}^{2}$ is the standard module on $\operatorname{SL}(2,\mathbf{Q})$. Such a classification may have interest in the problem of giving a definition of the notion ”good filtration” for torsion free finite rank comodules on $\operatorname{SL}(2,\mathbf{Z})$.
2. The adjoint representation of $\operatorname{GL}(2,A)$ and $\operatorname{SL}(2,A)$
Let $A$ be any commutative unital ring and let $x_{11},x_{12},x_{21},x_{22}$ be independent varaibles over $A$. Let $D:=x_{11}x_{22}-x_{12}x_{21}$ be the ”determinant” and let $F(x_{ij}):=D-1\in A[x_{ij}]$. Let $k[S]:=A[x_{ij}]/(F)$ and let $k[G]:=A[x_{ij},1/D]$. Let $\operatorname{SL}(2,A):=\operatorname{Spec}(k[S])$ and let $\operatorname{GL}(2,A):=\operatorname{Spec}(k[G])$. Let $\underline{A-alg}$ be the category of commutative $A$-algebras and let $\underline{Groups}$ be the category of groups. Define for $R=k[S],k[G]$ the functor
@@ 2. The adjoint representation of $\operatorname{GL}(2,A)$ and $\operatorname{SL}(2,A)$
by $h_{R}(B):=\operatorname{Hom}_{A-alg}(R,B)$. It follows $h_{R}(B)$ is the group of $2\times 2$ matrices $g:=(a_{ij})$ with coefficients $a_{ij}$ in $B$. If $R:=k[S]$ It follows $det(g)=1$ and if $R=k[G]$ it follows $det(g)$ is a unit in $B$. The affine schemes $\operatorname{GL}(2,A),\operatorname{SL}(2,A)$ are affine group scheme in the sense of the book of Demazure/Gabriel and Jantzen and the rings $k[G],k[S]$ are A-Hopf algebras. Define the following structure as a Hopf algebra on $k[G]$ and $k[S]$:
@@
In this section we define the adjoint representation of $\operatorname{GL}(2,A),\operatorname{SL}(2,A)$ on their Lie algebras using the functor of points $h_{G}(-)$.
@@
It follows $\operatorname{Lie}(G)^{F}(B)$ is the set of all $2\times 2$-matrices with coefficients in $B$. Hence $\operatorname{Lie}(G)^{F}(B)\cong\mathfrak{gl}(2,B)\cong\mathfrak{gl}(2,A)\otimes_{A}B$.
@@
$\mathfrak{gl}(2,A)_{a}(-):\underline{A-alg}\rightarrow\underline{Groups}$
@@
$\mathfrak{gl}(2,A)_{a}(B):=(\mathfrak{gl}(2,A)\otimes_{A}B,+),$
where $(\mathfrak{gl}(2,A)\otimes_{A}B,+)$ is the underlying abelian group of the $B$-module $\mathfrak{gl}(2,A)\otimes_{A}B$. For any map of $A$-algebras $\phi:B\rightarrow B_{1}$ we get an induced map
$\phi_{a}:\mathfrak{gl}(2,A)_{a}(B)\rightarrow\mathfrak{gl}(2,A)_{a}(B_{1})$
defined by $\phi_{a}(x\otimes b):=x\otimes\phi(b)$, hence the above definition gives a functor as claimed. There is an isomorphism of functors $\operatorname{Lie}(G)^{F}(-)\cong\mathfrak{gl}(2,A)_{a}(-)$.
We want to define an action of $h_{G}(-)$ on $\mathfrak{gl}(2,A)_{a}(-)$. Define for any $B\in\underline{A-alg}$ the following action:
$\sigma_{B}(-,-):h_{G}(B)\times\mathfrak{gl}(2,A)_{a}(B)\rightarrow\mathfrak{gl}(2,A)_{a}(B)$
@@
$\operatorname{Ad}^{F}:h_{G}(-)\rightarrow h_{\operatorname{GL}(\mathfrak{gl}(2,A))}(-)$
called the (left) adjoint representation of $G$ on $\mathfrak{gl}(2,A)$. There is similarly a right adjoint representation. We want to calculate the corresponding comodule map of the adjoint representation $\operatorname{Ad}^{F}$. We get a map
$\sigma_{k[G]}(-,-):h_{G}(k[G])\times\mathfrak{gl}(2,A)_{a}(k[G])\rightarrow\mathfrak{gl}(2,A)_{a}(k[G])$
@@
$\rho:\mathfrak{gl}(2,A)_{a}(k[G])\rightarrow\mathfrak{gl}(2,A)_{a}(k[G])$
@@
$\rho(z\otimes f):=\sigma(id_{k[G]},z\otimes f)\in\mathfrak{gl}(2,A)_{a}(k[G]):=\mathfrak{gl}(2,A)\otimes_{A}k[G].$
@@
$\operatorname{Ad}:\mathfrak{gl}(2,A)\rightarrow\mathfrak{gl}(2,A)\otimes_{A}k[G]$
@@
$\operatorname{Ad}(z):=\rho(z\otimes 1_{k[G]})\in gl(2,A)\otimes_{A}k[G].$
@@
Define the following matrices in $\mathfrak{gl}(2,A)$:
@@
It follows $\mathfrak{gl}(2,A)\cong A\{e_{ij}\}$ is a free $A$-module of rank $4$ on the basis $B:=\{e_{ij}\}$. We want to calculate the element $\operatorname{Ad}(e_{ij})$ for all $i,j$.
@@
You may easily check the above formulas define a comodule map $\operatorname{Ad}:\mathfrak{gl}(2,A)\rightarrow\mathfrak{gl}(2,A)\otimes_{A}k[G]$. We want to calculate the matrix corresponding to $\operatorname{Ad}$ in the basis $B$. The map $\operatorname{Ad}$ corresponds $1-1$ to a map (also denoted $\operatorname{Ad}$)
$\operatorname{Ad}:\mathfrak{gl}(2,A)\otimes_{A}k[G]\rightarrow\mathfrak{gl}(2,A)\otimes_{A}k[G].$
For any $C\in\underline{A-alg}$ we let the module $\mathfrak{gl}(2,A)\otimes_{A}C$ have the basis $B_{C}:=\{e_{ij}\otimes 1_{C}\}$. Here $1_{C}\in C$ is the multiplicative unit. We get the following matrix:
@@
$\operatorname{Ad}_{N}:\mathfrak{gl}(2,A)\otimes k[N]\rightarrow\mathfrak{gl}(2,A)\otimes k[N]$
@@
$\operatorname{Ad}_{T}:\mathfrak{gl}(2,A)\otimes k[T]\rightarrow\mathfrak{gl}(2,A)\otimes_{A}k[T]$
@@
Because of the isomorphim $\operatorname{Lie}(G)^{F}(-)\cong\mathfrak{gl}(2,A)_{a}(-)$ of functors there is an isomorphism $\operatorname{Lie}(\operatorname{GL}(2,A))\cong\mathfrak{gl}(2,A)$.
Let $\operatorname{Lie}(T):=A\{e_{11},e_{22}\}$ and $\operatorname{Lie}(\operatorname{GL}(2,A))_{R_{G}}:=A\{e_{12},e_{21}\}$. From the above calculation we get comodul structures
$\operatorname{Ad}_{R_{G}}:\operatorname{Lie}(\operatorname{GL}(2,A))_{R_{G}}\rightarrow\operatorname{Lie}(\operatorname{GL}(2,A))_{R_{G}}\otimes k[N]$
@@
and a direct sum of $k[N]$-comodules $\operatorname{Lie}(\operatorname{GL}(2,A))\cong\operatorname{Lie}(T)\oplus\operatorname{Lie}(\operatorname{GL}(2,A))_{R_{G}}$.
@@
$\operatorname{Lie}(\operatorname{GL}(2,A))\cong\operatorname{Lie}(T)\oplus\operatorname{Lie}(\operatorname{GL}(2,A))_{R_{G}}$
is the refined weight space decomposition of $\operatorname{Lie}(\operatorname{GL}(2,A))$. The set $R_{G}$ is the roots of $\operatorname{GL}(2,A)$.

### v4

# A refined Weyl character formula for comodules on $\operatorname{GL}_{2,A}$

@@ Abstract.
Let $A$ be any commutative unital ring and let $\operatorname{GL}_{2,A}$ be the general linear group on $A$ of rank $2$. I study the representation theory of $\operatorname{GL}_{2,A}$ and the symmetric powers $\operatorname{Sym}^{d}(V)$, where $(V,\Delta)$ is the standard right comodule on $\operatorname{GL}_{2,A}$. I prove a refined Weyl character formula for $\operatorname{Sym}^{d}(V)$. There is for any integer $d\geq 1$ a (canonical) refined weight space decomposition $\operatorname{Sym}^{d}(V)\cong\oplus_{i}\operatorname{Sym}^{d}(V)^{i}$ where each direct summand $\operatorname{Sym}^{d}(V)^{i}$ is a comodule on $N\subseteq\operatorname{GL}_{2,A}$. Here $N$ is the schematic normalizer of the diagonal torus $T\subseteq\operatorname{GL}_{2,A}$. I prove a character formula for the direct summands of $\operatorname{Sym}^{d}(V)$ for any integer $d\geq 1$. This refined Weyl character formula implies the classical Weyl character formula. As a Corollary I prove a refined Weyl character formula for the pull back $\operatorname{Sym}^{d}(V\otimes K)$ as a comodule on $\operatorname{GL}_{2,K}$ where $K$ is any field. I also prove that the direct summand $\operatorname{Sym}^{d}(V)^{i}$ is an irreducible $N$-comodule over any field. I moreover calculate explicit examples involving the symmetric powers, symmetric tensors and their duals.
@@ 1. Introduction
In a recent paper (see [9]) I studied the notion ”good filtration” for torsion free comodules on $\operatorname{SL}_{2,\mathbf{Z}}$ and I proved the existence of a ”virtual Clebsch-Gordan formula” for the tensor product $\operatorname{Sym}^{m}(V)\otimes\operatorname{Sym}^{n}(V)$ of the symmetric powers of the standard right comodule $(V,\Delta)$ on the group scheme $\operatorname{SL}_{2,A}$ where $A$ is an arbitrary commutative unital ring and $d\geq 1$ is an integer. This gives an infinite set of non trivial examples of torsion free finite rank comodules on $\operatorname{SL}_{2,\mathbf{Z}}$ equipped with a good filtration. In general one wants to give a definition of the notion ”good filtration” and to classify torsion free finite rank comodules with a good filtration. In [9] I found examples of non isomorphic comodules on $\operatorname{SL}_{2,\mathbf{Z}}$ that become isomorphic when we ”restrict to the generic fiber”. There is a canonical morphism of group schemes $\pi:\operatorname{SL}_{2,\mathbf{Q}}\rightarrow\operatorname{SL}_{2,\mathbf{Z}}$ and I proved existence of non isomorphic finite rank torsion free comodules $W_{1},W_{2}$ with $\pi^{*}W_{1}\cong\pi^{*}W_{2}$ an isomorphism of comodules on $\operatorname{SL}_{2,\mathbf{Q}}$. In order to define the notion ”good filtration” for comodules on $\operatorname{SL}_{2,\mathbf{Z}}$, we must for each integer $d\geq 1$ classify comodules $W$ on $\operatorname{SL}_{2,\mathbf{Z}}$ with the property that the pull back to $\operatorname{SL}_{2,\mathbf{Q}}$ becomes isomorphic to $\operatorname{Sym}^{d}(V)$ where $V:=\mathbf{Q}^{2}$ is the ”standard right comodule” of $\operatorname{SL}_{2,\mathbf{Q}}$. There is by [9] no highest weight theory or ”complete reducibility property” for torsion free comodules on $\operatorname{SL}_{2,\mathbf{Z}}$ and $\operatorname{GL}_{2,\mathbf{Z}}$, and for this reason we need new invariants and new methods in the study of such comodules and the aim of this paper is to introduce a new structure on a class of torsion free comodules on $\operatorname{SL}_{2,\mathbf{Z}}$.
In this paper I prove the following: Let $A$ be any commutative unital ring and let $G:=\operatorname{GL}_{2,A}$ or $\operatorname{SL}_{2,A}$. I define the schematic normalizer $N\subseteq G$, which is a closed subgroup scheme of $G$ containing the diagonal torus $T$ as a closed normal subgroup scheme. The schematic normalizer is isomorphic to the functor normalizer $N_{G}(T)$ as defined in [8]. Let $\sigma,q:N\times T\rightarrow N$ be the action and projection map.
@@
For any map $f:N\rightarrow Z$ to an affine scheme $Z$ with $f\circ\sigma=f\circ q$ there is a unique map $f^{*}:N/T\rightarrow Z$ with $f^{*}\circ\pi=f$. Hence the affine group scheme $N/T$ is uniquely determined up to isomorphism. If $G:=\operatorname{GL}_{2,A}$ it follows the sequence splits and there is an isomorphism of group schemes $N\cong T\times^{\sigma}N/T$ where the product on the right is the semi direct product.
@@
Then I prove the existence of a refined weight space decomposition of any symmetric power $W:=\operatorname{Sym}^{d}(V)$ where $d\geq 1$ is an integer and where $(V,\Delta)$ is the standard right comodule on $\operatorname{GL}_{2,A}$. There is a canonical decomposition $W\cong\oplus_{\lambda\in\Lambda}W_{\lambda}$ where each summand $W_{\lambda}$ is a comodule on the schematic normalizer $N\subseteq\operatorname{GL}_{2,A}$ of the diagonal torus $T\subseteq\operatorname{GL}_{2,A}$. Then I prove a refined Weyl character formula for each summand $W_{\lambda}$ of $W$. The refined character formula implies the classical Weyl character formula (this is Theorem 6.4 in the paper):
@@
In Example 6.6 we relate the refined weight space decomposition to irreducible finite dimensional modules over a field $k$ of positive charcacteristic. If $G:=\operatorname{SL}_{n,k}$ and if $S_{\lambda}(V)$ is a Weyl module, there is a unique finite dimensional irreducible module $V(\lambda)\subsetneq S_{\lambda}(V)$. The irreducible module $V(\lambda)$ has a relation with the refined weight space decomposition for $S_{\lambda}(V)$.
In Theorem 7.1 I prove that the direct summand $\operatorname{Sym}^{d}(V)^{i}$ is an irreducible module on $N$ over any field $k$.
As an application of the methods introduced in the paper I prove the following Theorems (see Theorem 8.3). Let $G:=\operatorname{GL}_{2,A}$ where $A$ is any commutative ring and let $(V,\Delta)$ be the standard left comodule on $G$. Let $\operatorname{Sym}^{2}(V),\operatorname{S}_{2}(V),\operatorname{Sym}^{2}(V)^{*}$ and $\operatorname{S}_{2}(V)^{*}$ be the symmetric power, the symmetric tensor and the dual of the symmetric power and symmetric tensor of $V$. The symmetric tensor is the invariants $\operatorname{S}_{2}(V)\subseteq V\otimes_{A}V$ under the canonical action of the symmetric group on two elements.
@@
Let $G:=\operatorname{GL}_{2,A}$ with $A$ any commutative ring, and let $N\subseteq G$ be the schematic normalizer of the diagonal torus. There is an isomorphism of $k[N]$-comodules (and $k[T]$-comodules) $\operatorname{Sym}^{2}(V)\cong\operatorname{S}_{2}(V)$. Let $G:=\operatorname{GL}_{2,\mathbf{Z}}$ with $\mathbf{Z}$ the ring of integers. It follows the $k[G]$-comodules $\operatorname{Sym}^{2}(V)$ and $\operatorname{S}_{2}(V)$ are not isomorphic.
@@
It is hoped the results in this paper may be interesting in the study of torsion free comodules on $\operatorname{SL}_{2,A}$ and $\operatorname{GL}_{2,A}$ in the future. One problem one wants to study is the problem of classifying $\operatorname{SL}_{2,\mathbf{Z}}$ modules $W$ whose base change to $\operatorname{SL}_{2,\mathbf{Q}}$ are isomorphic to $\operatorname{Sym}^{d}(\mathbf{Q}^{2})$ where $\mathbf{Q}^{2}$ is the standard module on $\operatorname{SL}_{2,\mathbf{Q}}$. Such a classification may have interest in the problem of giving a definition of the notion ”good filtration” for torsion free finite rank comodules on $\operatorname{SL}_{2,\mathbf{Z}}$. Such a notion may have applications in the study of the representation theory of $\operatorname{SL}_{n,k}$ and $\operatorname{GL}_{n,k}$ over an arbitrary field $k$.
2. The adjoint representation of $\operatorname{GL}_{2,A}$ and $\operatorname{SL}_{2,A}$
Let $A$ be any commutative unital ring and let $x_{11},x_{12},x_{21},x_{22}$ be independent varaibles over $A$. Let $D:=x_{11}x_{22}-x_{12}x_{21}$ be the ”determinant” and let $F(x_{ij}):=D-1\in A[x_{ij}]$. Let $k[S]:=A[x_{ij}]/(F)$ and let $k[G]:=A[x_{ij},1/D]$. Let $\operatorname{SL}_{2,A}:=\operatorname{Spec}(k[S])$ and let $\operatorname{GL}_{2,A}:=\operatorname{Spec}(k[G])$. Let $\underline{A-alg}$ be the category of commutative $A$-algebras and let $\underline{Groups}$ be the category of groups. Define for $R=k[S],k[G]$ the functor
@@ 2. The adjoint representation of $\operatorname{GL}_{2,A}$ and $\operatorname{SL}_{2,A}$
by $h_{R}(B):=\operatorname{Hom}_{A-alg}(R,B)$. It follows $h_{R}(B)$ is the group of $2\times 2$ matrices $g:=(a_{ij})$ with coefficients $a_{ij}$ in $B$. If $R:=k[S]$ It follows $det(g)=1$ and if $R=k[G]$ it follows $det(g)$ is a unit in $B$. The affine schemes $\operatorname{GL}_{2,A},\operatorname{SL}_{2,A}$ are affine group scheme in the sense of the book of Demazure/Gabriel and Jantzen and the rings $k[G],k[S]$ are A-Hopf algebras. Define the following structure as a Hopf algebra on $k[G]$ and $k[S]$:
@@
In this section we define the adjoint representation of $\operatorname{GL}_{2,A},\operatorname{SL}_{2,A}$ on their Lie algebras using the functor of points $h_{G}(-)$.
@@
It follows $\operatorname{Lie}(G)^{F}(B)$ is the set of all $2\times 2$-matrices with coefficients in $B$. Hence $\operatorname{Lie}(G)^{F}(B)\cong\mathfrak{gl}_{2,B}\cong\mathfrak{gl}_{2,A}\otimes_{A}B$.
@@
$(\mathfrak{gl}_{2,A})_{a}(-):\underline{A-alg}\rightarrow\underline{Groups}$
@@
$(\mathfrak{gl}_{2,A})_{a}(B):=(\mathfrak{gl}_{2,A}\otimes_{A}B,+),$
where $(\mathfrak{gl}_{2,A}\otimes_{A}B,+)$ is the underlying abelian group of the $B$-module $\mathfrak{gl}_{2,A}\otimes_{A}B$. For any map of $A$-algebras $\phi:B\rightarrow B_{1}$ we get an induced map
$\phi_{a}:(\mathfrak{gl}_{2,A})_{a}(B)\rightarrow(\mathfrak{gl}_{2,A})_{a}(B_{1})$
defined by $\phi_{a}(x\otimes b):=x\otimes\phi(b)$, hence the above definition gives a functor as claimed. There is an isomorphism of functors $\operatorname{Lie}(G)^{F}(-)\cong(\mathfrak{gl}_{2,A})_{a}(-)$.
We want to define an action of $h_{G}(-)$ on $(\mathfrak{gl}_{2,A})_{a}(-)$. Define for any $B\in\underline{A-alg}$ the following action:
$\sigma_{B}(-,-):h_{G}(B)\times(\mathfrak{gl}_{2,A})_{a}(B)\rightarrow(\mathfrak{gl}_{2,A})_{a}(B)$
@@
$\operatorname{Ad}^{F}:h_{G}(-)\rightarrow h_{\operatorname{GL}(\mathfrak{gl}_{2,A})}(-)$
called the (left) adjoint representation of $G$ on $\mathfrak{gl}_{2,A}$. There is similarly a right adjoint representation. We want to calculate the corresponding comodule map of the adjoint representation $\operatorname{Ad}^{F}$. We get a map
$\sigma_{k[G]}(-,-):h_{G}(k[G])\times(\mathfrak{gl}_{2,A})_{a}(k[G])\rightarrow(\mathfrak{gl}_{2,A})_{a}(k[G])$
@@
$\rho:(\mathfrak{gl}_{2,A})_{a}(k[G])\rightarrow(\mathfrak{gl}_{2,A})_{a}(k[G])$
@@
$\rho(z\otimes f):=\sigma(id_{k[G]},z\otimes f)\in(\mathfrak{gl}_{2,A})_{a}(k[G]):=\mathfrak{gl}_{2,A}\otimes_{A}k[G].$
@@
$\operatorname{Ad}:\mathfrak{gl}_{2,A}\rightarrow\mathfrak{gl}_{2,A}\otimes_{A}k[G]$
@@
$\operatorname{Ad}(z):=\rho(z\otimes 1_{k[G]})\in\mathfrak{gl}_{2,A}\otimes_{A}k[G].$
@@
Define the following matrices in $\mathfrak{gl}_{2,A}$:
@@
It follows $\mathfrak{gl}_{2,A}\cong A\{e_{ij}\}$ is a free $A$-module of rank $4$ on the basis $B:=\{e_{ij}\}$. We want to calculate the element $\operatorname{Ad}(e_{ij})$ for all $i,j$.
@@
You may easily check the above formulas define a comodule map $\operatorname{Ad}:\mathfrak{gl}_{2,A}\rightarrow\mathfrak{gl}_{2,A}\otimes_{A}k[G]$. We want to calculate the matrix corresponding to $\operatorname{Ad}$ in the basis $B$. The map $\operatorname{Ad}$ corresponds $1-1$ to a map (also denoted $\operatorname{Ad}$)
$\operatorname{Ad}:\mathfrak{gl}_{2,A}\otimes_{A}k[G]\rightarrow\mathfrak{gl}_{2,A}\otimes_{A}k[G].$
For any $C\in\underline{A-alg}$ we let the module $\mathfrak{gl}_{2,A}\otimes_{A}C$ have the basis $B_{C}:=\{e_{ij}\otimes 1_{C}\}$. Here $1_{C}\in C$ is the multiplicative unit. We get the following matrix:
@@
$\operatorname{Ad}_{N}:\mathfrak{gl}_{2,A}\otimes k[N]\rightarrow\mathfrak{gl}_{2,A}\otimes k[N]$
@@
$\operatorname{Ad}_{T}:\mathfrak{gl}_{2,A}\otimes k[T]\rightarrow\mathfrak{gl}_{2,A}\otimes_{A}k[T]$
@@
Because of the isomorphim $\operatorname{Lie}(G)^{F}(-)\cong(\mathfrak{gl}_{2,A})_{a}(-)$ of functors there is an isomorphism $\operatorname{Lie}(\operatorname{GL}_{2,A})\cong\mathfrak{gl}_{2,A}$.
Let $\operatorname{Lie}(T):=A\{e_{11},e_{22}\}$ and $\operatorname{Lie}(\operatorname{GL}_{2,A})_{R_{G}}:=A\{e_{12},e_{21}\}$. From the above calculation we get comodul structures
$\operatorname{Ad}_{R_{G}}:\operatorname{Lie}(\operatorname{GL}_{2,A})_{R_{G}}\rightarrow\operatorname{Lie}(\operatorname{GL}_{2,A})_{R_{G}}\otimes k[N]$
@@
and a direct sum of $k[N]$-comodules $\operatorname{Lie}(\operatorname{GL}_{2,A})\cong\operatorname{Lie}(T)\oplus\operatorname{Lie}(\operatorname{GL}_{2,A})_{R_{G}}$.
@@
$\operatorname{Lie}(\operatorname{GL}_{2,A})\cong\operatorname{Lie}(T)\oplus\operatorname{Lie}(\operatorname{GL}_{2,A})_{R_{G}}$
is the refined weight space decomposition of $\operatorname{Lie}(\operatorname{GL}_{2,A})$. The set $R_{G}$ is the roots of $\operatorname{GL}_{2,A}$.